# Finite Distance Corrections to Vacuum Birefringence in Strong Gravitational and Electromagnetic Fields

### v1

@@ IV.3 The finite-distance correction series
The expression derived in Eq. (28) provides the exact first-order differential birefringence angle. However, to make the result useful for observational tests (e.g., with magnetars), we must separate the result into the standard asymptotic value (observable at infinity) and the finite-distance corrections that arise when the observer or source is near the lens.
@@
Expanding the integral in Eq. (28) around these limits, we isolate the terms that depend on $u_{\text{R}}$ and $u_{\text{S}}$. The total differential deflection $\Delta\alpha$ can be written as:
$\Delta\alpha(r_{\text{R}},r_{\text{S}})=\Delta\alpha_{\infty}-\left(\delta\alpha(r_{\text{R}})+\delta\alpha(r_{\text{S}})\right),$ (29)
where $\Delta\alpha_{\infty}$ is the asymptotic birefringence obtained by integrating from 0 to $\pi$, and $\delta\alpha(r)$ represents the reduction in the accumulated phase shift due to the finite path length.
@@
$\delta\alpha(r_{\text{R,S}})\approx\int_{0}^{\arcsin(b/r_{\text{R,S}})}\frac{C_{n}}{(b/\sin\phi)^{n}}d\phi.$ (30)
For physically relevant models like vacuum polarization (QED) or Born-Infeld theory, the effective curvature difference decays rapidly with distance. Consequently, the finite-distance correction can be expressed as a series expansion in terms of the ratio of the impact parameter to the observer distance. The general form of the finite-distance birefringence is:
$\boxed{\Delta\alpha(r_{\text{R}},r_{\text{S}})=\Delta\alpha_{\infty}\left[1-c_{\text{R}}\left(\frac{b}{r_{\text{R}}}\right)^{n}-c_{\text{S}}\left(\frac{b}{r_{\text{S}}}\right)^{n}+\dots\right],}$ (31)
where the exponent $n$ and the coefficients $c_{\text{R,S}}$ are determined by the asymptotic falloff of the effective optical curvature $K^{(1)}$ and the background field configuration. For example, in a background driven by a magnetic monopole, the curvature scales differently than in a dipole field, altering the leading-order distance correction. We will derive the explicit values for $n$ for the Euler-Heisenberg and Born-Infeld models in the subsequent section.
@@ V Applications to physical metrics
Having established the general perturbative framework for finite-distance birefringence in Section IV, we now apply this formalism to specific physical scenarios of interest. The power of the expression derived in Eq. (28) lies in its universality; it allows us to compute the observable differential deflection for any non-linear electrodynamics (NLED) theory, provided the background metric and the specific form of the Lagrangian coupling are known. In this section, we calculate the explicit birefringence corrections for two distinct cases: the Euler-Heisenberg effective action of Quantum Electrodynamics (QED) and the Born-Infeld non-linear extension of Maxwell’s theory. For each case, we isolate the finite-distance corrections described by the series expansion in Eq. (31) , determining the model-dependent coefficients that could be probed by next-generation X-ray polarimetry.
@@ V.1.1 The effective Lagrangian and metric perturbations
$\mathcal{L}_{\text{EH}}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{\beta_{\text{QED}}}{4}\left[(F_{\mu\nu}F^{\mu\nu})^{2}+\frac{7}{4}(F_{\mu\nu}\tilde{F}^{\mu\nu})^{2}\right],$ (32)
where the coupling parameter is $\beta_{\text{QED}}=\frac{2\alpha_{\text{EM}}^{2}}{45m_{e}^{4}}$, with $m_{e}$ representing the electron mass. Such a form matches the perturbative structure assumed in Eq. (21), where $\beta_{\text{QED}}$ serves as the expansion parameter.
For the background field configuration, we assume a static, spherically symmetric magnetic monopole solution with magnetic charge $Q_{m}$ in a Schwarzschild spacetime. While magnetic monopoles remain hypothetical, this configuration is standard in theoretical birefringence studies because it preserves spherical symmetry, greatly simplifying the evaluation of the optical metric functions $A_{\pm}$ and $B_{\pm}$. The background electromagnetic field tensor has a single non-vanishing component $F_{\theta\phi}=Q_{m}\sin\theta$, corresponding to a radial magnetic field $|B|\approx Q_{m}/r^{2}$.
In this configuration, the field invariant is $F_{\mu\nu}F^{\mu\nu}=2Q_{m}^{2}/r^{4}$ (and $F\tilde{F}=0$). Substituting this into the effective metric definitions from the previous section, the perturbation functions $\mathcal{A}_{\pm}$, $\mathcal{B}_{\pm}$, and $\mathcal{H}_{\pm}$ in Eq. (21) take the form:
$\mathcal{A}_{\pm}(r)=\mathcal{B}_{\pm}(r)=\lambda_{\pm}\frac{Q_{m}^{2}}{r^{4}},\quad\mathcal{H}_{\pm}(r)=0,$ (33)
where $\lambda_{\pm}$ are numerical coefficients determined by the polarization mode (typically 14/45 and 8/45 for the two modes in QED). Note that for a pure magnetic background, the angular deformation $\mathcal{H}$ vanishes, simplifying the orbit equation derived in Eq. (23).
@@ V.1.2 The finite-distance birefringence profile
To compute the differential bending angle $\Delta\alpha$, we utilize the decomposition into refractive curvature and path-induced contributions derived in Eq. (28). The difference in the effective optical Gaussian curvature $\Delta K$ is the dominant source of birefringence. Given the $1/r^{4}$ scaling of the metric perturbations in Eq. (33), the effective curvature difference scales as the second derivative of the metric perturbation:
$\Delta K(r)\approx\frac{d^{2}}{dr^{2}}\left(\beta_{\text{QED}}\Delta\lambda\frac{Q_{m}^{2}}{r^{4}}\right)\sim\beta_{\text{QED}}\frac{Q_{m}^{2}}{r^{6}}.$ (34)
Inserting Eq. (34) into the refractive curvature integral of Eq. (28), we integrate over the domain bounded by the photon path $r(\phi)\approx b/\sin\phi$. The area element contributes a factor of $\sqrt{\gamma}\sim r$, making the integrand scale as $\Delta(K\sqrt{\gamma})\sim r^{-5}$.
Performing the integration over the radial coordinate first, as prescribed in Eq. (31), we find that the accumulated birefringence falls off rapidly with the impact parameter $b$. Specifically, the asymptotic birefringence scales as $\Delta\alpha_{\infty}\propto\beta_{\text{QED}}Q_{m}^{2}/b^{4}$. The steep fall-off ($n=4$) reflects the rapid decay of the dipole vacuum polarization cloud surrounding the central object.
We can now explicitly write the finite-distance correction for QED vacuum birefringence. Substituting the scaling $n=4$ into the general series solution Eq. (31), we obtain:
$\boxed{\Delta\alpha_{\text{QED}}(r_{\text{R}},r_{\text{S}})=\Delta\alpha_{\infty}\left[1-c_{\text{R}}\left(\frac{b}{r_{\text{R}}}\right)^{4}-c_{\text{S}}\left(\frac{b}{r_{\text{S}}}\right)^{4}+\mathcal{O}\left(\frac{b^{6}}{r^{6}}\right)\right].}$ (35)
@@ V.1.3 Physical implications
Equation (35) highlights a critical constraint for observational campaigns. The correction term scales as the fourth power of the ratio $b/r$. Unlike the gravitational deflection of light (which scales as $b/r$ in the finite-distance correction), the QED birefringence is highly localized near the source. Consequently, for an observer at infinity ($r_{\text{R}}\to\infty$), the finite-distance correction vanishes rapidly. However, if the source is located within the strong-field region (small $r_{\text{S}}$, e.g., emission from the surface of a magnetar), the truncation term $c_{\text{S}}(b/r_{\text{S}})^{4}$ becomes significant. The implication is that the standard asymptotic formulas over-estimate the expected polarization rotation for surface emission models, necessitating the use of Eq. (35) for precise polarimetric modeling of magnetars.
@@ V.2 Case Study 2: The Born-Infeld black Hole
While Quantum Electrodynamics provides the standard model for vacuum birefringence, Born-Infeld (BI) electrodynamics offers a classical non-linear extension of Maxwell’s theory, originally proposed to regularize the self-energy of point charges. In the context of strong-field gravity, BI theory often serves as a prototype for effective actions arising in string theory. Unlike the perturbative QED action, which is an approximation valid only for fields weak compared to the Schwinger limit, the Born-Infeld action is defined to all orders, though we treat it here within the perturbative framework established in Section IV to allow for a direct comparison of the finite-distance effects.
@@ V.2.1 Lagrangian and weak-field limit
The Born-Infeld Lagrangian density is given by:
$\mathcal{L}_{\text{BI}}=\gamma^{2}\left(1-\sqrt{1+\frac{1}{2\gamma^{2}}F_{\mu\nu}F^{\mu\nu}-\frac{1}{16\gamma^{4}}(F_{\mu\nu}\tilde{F}^{\mu\nu})^{2}}\right),$ (36)
@@
where $\gamma$ is the Born-Infeld coupling parameter (having dimensions of field strength), which sets the scale at which non-linearities become dominant. In the limit $\gamma\to\infty$, we recover Maxwell’s theory. To apply the weak-field expansion formalism of Eq. (21), we identify the perturbative parameter $\beta_{\text{BI}}\propto 1/\gamma^{2}$.
@@
Expanding Eq. (36) for weak fields ($|F|\ll\gamma$) and retaining terms up to the first order in the non-linearity (order $F^{4}$), we obtain:
$\mathcal{L}_{\text{BI}}\approx-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{1}{32\gamma^{2}}(F_{\mu\nu}F^{\mu\nu})^{2}+\mathcal{O}\left(\frac{1}{\gamma^{4}}\right).$ (37)
The above expansion is valid in the region $r\gg r_{core}$ where the field strength is significantly below the critical Born-Infeld scale $\gamma$. We assume the source and observer are located within this perturbative domain, avoiding the non-analytic behavior near the field singularity. Comparing this expansion with the QED Lagrangian in Eq. (32), we observe a crucial structural similarity: both theories introduce a quartic correction to the field invariant. However, a key distinction arises in the polarization structure. The Euler-Heisenberg action includes the $(F_{\mu\nu}\tilde{F}^{\mu\nu})^{2}$ term (associated with parity-odd effects or magnetic-electric mixing), which is absent in the Born-Infeld expansion at this order for purely magnetic or purely electric backgrounds where $F\tilde{F}=0$.
For the magnetic monopole background considered in the previous subsection ($F_{\mu\nu}F^{\mu\nu}=2Q_{m}^{2}/r^{4}$), the metric perturbation functions $\mathcal{A}_{\pm}$ and $\mathcal{B}_{\pm}$ are determined by the coefficient of the quartic term. For Born-Infeld, the coupling is strictly positive, but the resulting birefringence usually exhibits a sign difference relative to QED depending on the specific mode definitions.
V.2.2 Finite-distance correction coefficients
We proceed to calculate the finite-distance birefringence using the master equation, Eq. (28). The refractive curvature difference $\Delta K$ is again driven by the second radial derivative of the metric perturbation. Since the leading-order correction in Eq. (37) scales as $(F^{2})^{2}\sim r^{-8}$ in the Lagrangian, the metric perturbation scales as $1/r^{4}$ (similar to QED).
Consequently, the effective optical curvature difference $\Delta K_{\text{BI}}$ follows the same radial power law as the QED case:
$\Delta K_{\text{BI}}(r)\sim\frac{1}{\gamma^{2}}\frac{Q_{m}^{2}}{r^{6}}.$ (38)
Following the integration procedure outlined in Eq. (30) and (31), the finite-distance correction is dominated by the same $n=4$ decay index found in the Euler-Heisenberg analysis, which is considered as a univers consequence of the identical power-counting of the leading-order non-linearity in 4 dimensions.
However, the magnitude of the effect differs. The finite-distance birefringence for the Born-Infeld monopole is:
$\boxed{\Delta\alpha_{\text{BI}}(r_{\text{R}},r_{\text{S}})=\frac{\Delta\alpha_{\infty}^{\text{BI}}}{Q_{m}^{2}}\left[1-\tilde{c}_{\text{R}}\left(\frac{b}{r_{\text{R}}}\right)^{4}-\tilde{c}_{\text{S}}\left(\frac{b}{r_{\text{S}}}\right)^{4}\right],}$ (39)
where the asymptotic value $\Delta\alpha_{\infty}^{\text{BI}}$ is directly proportional to the ratio $Q_{m}^{2}/(\gamma^{2}b^{4})$.
V.2.3 Comparison of QED and Born-Infeld signatures
While the radial scaling ($n=4$) suggests that disentangling QED from Born-Infeld effects based solely on distance dependence is difficult for a monopole source, the coefficients $\tilde{c}_{\text{R,S}}$ provide a potential discriminator if the impact parameter $b$ can be independently constrained.
Furthermore, if we were to relax the assumption of a monopole background and consider a physical magnetic dipole (typical of pulsars), the radial scaling would change. For a dipole, the background field falls off as $B\sim r^{-3}$, implying $F^{2}\sim r^{-6}$. The resulting metric perturbation would scale as $r^{-6}$, and the curvature difference as $\Delta K\sim r^{-8}$. Inserting this into the integral of Eq. (28), the finite-distance correction would then follow a steeper power law:
$\delta\alpha_{\text{dipole}}(r)\sim\int^{\arcsin(b/r)}(\sin\phi)^{6}d\phi\implies n=6,$ (40)
which demonstrates that the finite-distance correction series in Eq. (31) is sensitive to the multipole structure of the background field. For magnetars, where the field structure may contain higher-order multipoles near the surface, the truncation terms $c_{\text{S}}(b/r_{\text{S}})^{n}$ will be dominated by the highest multipole component, making the finite-distance correction even more critical than the monopole approximation suggests.
Figure 1: The normalized differential deflection angle $\Delta\alpha/\Delta\alpha_{\infty}$ as a function of the source distance $r_{\text{S}}$ scaled by the impact parameter $b$. The solid black line corresponds to the magnetar dipole model ($n=6$), while the dashed red line represents the magnetic monopole reference ($n=4$). The curves demonstrate that the dipole signal converges more rapidly to the asymptotic limit, indicating that the birefringent effect is more tightly confined to the strong-field region. For sources at $r_{\text{S}}\gtrsim 4b$, the finite-distance suppression becomes negligible ($5\%$)
The decomposition of the finite-distance corrections reveals a fundamental feature of vacuum birefringence in compact object environments: strong spatial localization. As shown in Fig. 1, the normalized birefringence signal for the physical magnetar model (dipole scaling, $n=6$) converges to the asymptotic limit significantly faster than the monopole reference ($n=4$). For a source located at a distance $r_{\text{S}}\approx 4b$ (where $b$ is the impact parameter), the observable phase shift has already captured $95\%$ of the total theoretical value, which implies that for binary systems where the X-ray source is a companion star located at $r_{\text{S}}\gg R_{\text{NS}}$, the finite-distance corrections are negligible, and the standard asymptotic scattering formulas are valid.

### v2

@@ IV.3 The finite-distance correction series
The expression derived in Eq. (IV.8) provides the exact first-order differential birefringence angle. However, to make the result useful for observational tests (e.g., with magnetars), we must separate the result into the standard asymptotic value (observable at infinity) and the finite-distance corrections that arise when the observer or source is near the lens.
@@
Expanding the integral in Eq. (IV.8) around these limits, we isolate the terms that depend on $u_{\text{R}}$ and $u_{\text{S}}$. The total differential deflection $\Delta\alpha$ can be written as:
$\Delta\alpha(r_{\text{R}},r_{\text{S}})=\Delta\alpha_{\infty}-\left(\delta\alpha(r_{\text{R}})+\delta\alpha(r_{\text{S}})\right),$ (IV.10)
where $\Delta\alpha_{\infty}$ denotes the two-sided asymptotic scattering value obtained in the limit $r_{S},r_{R}\to\infty$ (so that $\phi_{S}\to 0$ and $\phi_{R}\to\pi$ for fixed impact parameter $b$), i.e. by extending the integration over the full in-and-out trajectory used in the Ishihara-Ono-Asada-Suzuki construction.
@@
$\delta\alpha(r_{\text{R,S}})\approx\int_{0}^{\arcsin(b/r_{\text{R,S}})}\frac{C_{n}}{(b/\sin\phi)^{n}}d\phi.$ (IV.11)
For physically relevant birefringent models such as vacuum polarization (Euler-Heisenberg QED), the effective curvature difference $\Delta(K\sqrt{\gamma})$ decays rapidly with distance. Consequently, the finite-distance correction can be expressed as a series expansion in terms of the ratio of the impact parameter to the observer distance. The general form of the finite-distance birefringence is:
$\Delta\alpha(r_{\text{R}},r_{\text{S}})=\Delta\alpha_{\infty}\left[1-c_{\text{R}}\left(\frac{b}{r_{\text{R}}}\right)^{n}-c_{\text{S}}\left(\frac{b}{r_{\text{S}}}\right)^{n}+\dots\right],$ (IV.12)
where the exponent $n$ is fixed by the leading large-$r$ scaling of the effective curvature difference $\Delta(K\sqrt{\gamma})\sim r^{-n}$ (hence $n$ is explicit once the far-field decay of the background field and the NLED model are specified). By contrast, $c_{S}$ and $c_{R}$ are dimensionless $\mathcal{O}(1)$ coefficients that encode the normalization of $\Delta\alpha_{\infty}$ (two-sided scattering) and the relative weight of the truncated asymptotic tail compared to the non-asymptotic part of the curvature flux; they are therefore not fixed by power counting alone. In practice, for a given model and geometry, $c_{S}$ and $c_{R}$ are obtained by matching the asymptotic expansion (IV.12) to the exact first-order result (IV.8) (numerically if needed). In Sec. V we compute $n$ explicitly for the Euler-Heisenberg monopole benchmark and quote benchmark values for $c_{S,R}$ in the surface-emission configuration.
IV.4 Domain of validity and expected accuracy
Our analytic results combine three controlled approximations, and are intended to be used only within their overlap regime: (i) For the weak-NLED expansion, the perturbative mapping in Sec. IV assumes that the polarization-dependent metric deformations are small,
$|\beta\,a_{\pm}(r)|\ll 1,\qquad|\beta\,b_{\pm}(r)|\ll 1,\qquad|\beta\,h_{\pm}(r)|\ll 1,$ (IV.13)
equivalently that the relevant dimensionless combinations built from the Lagrangian derivatives (e.g. $\beta F$, $\beta G$) remain perturbative along the photon path. For Euler-Heisenberg QED in a purely magnetic background, this corresponds to the standard condition $\epsilon_{\rm QED}\sim\frac{\alpha_{\rm EM}}{45\pi}\left(\frac{\mathcal{B}}{B_{Q}}\right)^{2}\ll 1$ (with $B_{Q}=m_{e}^{2}/e$), so that omitted higher-order terms scale as $\mathcal{O}(\epsilon_{\rm QED}^{2})$. (ii) In evaluating the finite-distance coefficients we use the background orbit $u_{0}(\phi)=\sin\phi/b$ and the associated limit expansions for $\phi_{S,R}$, which corresponds to retaining the leading weak-deflection geometry. Corrections from using the exact Schwarzschild background orbit enter parametrically at $\mathcal{O}(M/b)$ in the geometric limits and at higher order in the curvature flux; thus the leading truncation exponents $n$ are robust, while the order-unity coefficients $c_{S},c_{R}$ inherit $\mathcal{O}(M/b)$ uncertainties in the weak-field treatment. (iii) Throughout we place the electromagnetic configuration on a fixed background geometry (e.g. Schwarzschild) and neglect electromagnetic backreaction on the spacetime metric. This is appropriate when the electromagnetic stress-energy is small compared to the curvature scale set by the compact object (e.g. for neutron stars, $\mathcal{B}^{2}/8\pi\ll\rho c^{2}$), so that the dominant gravitational lensing is governed by the stellar mass rather than by the field energy.
For representative magnetar surface fields $\mathcal{B}_{\rm surf}\sim 10^{14}--10^{15}\,\mathrm{G}$, one has $\epsilon_{\rm QED}\lesssim 10^{-2}$-$10^{-1}$, so the leading-order Euler-Heisenberg birefringent corrections retained here are expected to be accurate at the few-percent level (with higher-order QED corrections becoming important only for substantially larger $B$). For gravitational compactness $M/R_{\rm NS}\sim 0.2$, the weak-deflection approximation is not extremely small for near-surface rays, so precise inference of order-unity coefficients should ultimately use the exact background orbit; the finite-distance scaling and geometric truncation mechanism derived in Secs. III-IV remain the central qualitative prediction.
@@ V Applications to physical metrics
Having established the general perturbative framework for finite-distance birefringence in Section IV, we now apply this formalism to specific physical scenarios of interest. The power of the expression derived in Eq. (IV.8) lies in its universality; it allows us to compute the observable differential deflection for any NLED theory, provided the background metric and the specific form of the Lagrangian coupling are known. We stress that while the decomposition in Eq. (IV.8) into a bulk term and a boundary (domain-shift) term depends on the chosen reference orbit used to linearize the domains, the sum equals the direct difference of Gauss-Bonnet integrals and is therefore the invariant observable $\Delta\alpha$ to $\mathcal{O}(\beta)$. In this section, we compute the explicit finite-distance birefringence for the Euler-Heisenberg effective action of QED [40, 4], and we then revisit Born-Infeld electrodynamics as a non-birefringent consistency check. With our observable defined as $\Delta\alpha\equiv\alpha_{+}-\alpha_{-}$, Born-Infeld theory gives $\Delta\alpha_{\rm BI}=0$ identically [13, 67], while any BI modification of the individual bending angles is polarization independent and therefore lies outside $\Delta\alpha$.
@@ V.1.1 The effective Lagrangian and metric perturbations
$\mathcal{L}_{\text{EH}}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{\beta_{\text{QED}}}{4}\left[(F_{\mu\nu}F^{\mu\nu})^{2}+\frac{7}{4}(F_{\mu\nu}\tilde{F}^{\mu\nu})^{2}\right],$ (V.1)
where the coupling parameter is $\beta_{\text{QED}}=\frac{2\alpha_{\text{EM}}^{2}}{45m_{e}^{4}}$, with $m_{e}$ representing the electron mass. Such a form matches the perturbative structure assumed in Eq. (IV.1), where $\beta_{\text{QED}}$ serves as the expansion parameter.
For the background field configuration, we assume a static, spherically symmetric magnetic monopole solution with magnetic charge $Q_{m}$ in a Schwarzschild spacetime. While magnetic monopoles remain hypothetical, this configuration is standard in theoretical birefringence studies because it preserves spherical symmetry, greatly simplifying the evaluation of the optical metric functions $A_{\pm}$ and $B_{\pm}$. The background electromagnetic field tensor has a single non-vanishing component $F_{\theta\phi}=Q_{m}\sin\theta$, corresponding to a radial magnetic field $|\mathcal{B}(r)|\simeq Q_{m}/r^{2}$.
In this configuration one has
$F_{\mu\nu}F^{\mu\nu}=\frac{2Q_{m}^{2}}{r^{4}},\qquad F_{\mu\nu}\tilde{F}^{\mu\nu}=0,$ (V.2)
i.e. in our invariant conventions (Sec. II) $F=\tfrac{1}{4}F_{\mu\nu}F^{\mu\nu}=Q_{m}^{2}/(2r^{4})$ and $G=0$.
For later use it is convenient to rewrite Eq. (V.1) in terms of $(F,G)$:
$L_{\rm EH}(F,G)=-F+\beta_{\rm QED}\bigl(4F^{2}+7G^{2}\bigr).$ (V.3)
Hence
$L_{F}=-1+8\beta_{\rm QED}F,\qquad L_{FF}=8\beta_{\rm QED},\qquad L_{GG}=14\beta_{\rm QED},\qquad L_{FG}=0,$ (V.4)
evaluated on the background ($G=0$).
For the monopole background, the only non-vanishing components of the contraction $T^{\mu\nu}\equiv F^{\mu}{}_{\lambda}F^{\nu\lambda}$ are angular. Explicitly, with $F_{\theta\phi}=Q_{m}\sin\theta$ one finds
$T^{tt}=T^{rr}=0,\qquad T^{\theta\theta}=\frac{Q_{m}^{2}}{r^{6}},\qquad T^{\phi\phi}=\frac{Q_{m}^{2}}{r^{6}\sin^{2}\theta}.$ (V.5)
Substituting into the effective-metric definition (Sec. II),
$g^{\mu\nu}_{{\rm eff},\pm}=L_{F}\,g^{\mu\nu}-4\Lambda_{\pm}\,T^{\mu\nu},$ (V.6)
shows that (up to the common conformal factor $L_{F}$) the polarization-dependent corrections enter precisely through the angular components. In particular,
$\frac{g^{rr}_{{\rm eff},\pm}}{g^{rr}}=L_{F},\qquad\frac{g^{\theta\theta}_{{\rm eff},\pm}}{g^{\theta\theta}}=L_{F}-4\,\Lambda_{\pm}\frac{Q_{m}^{2}}{r^{4}},$ (V.7)
and therefore, using the definition of $H_{\pm}$ (Sec. II),
$H_{\pm}(r)=\frac{g^{\theta\theta}_{{\rm eff}\,\pm}/g^{\theta\theta}}{g^{rr}_{{\rm eff}\,\pm}/g^{rr}}=1-\frac{4\Lambda_{\pm}}{L_{F}}\frac{Q_{m}^{2}}{r^{4}}=1+\beta_{\rm QED}\,h_{\pm}(r)+\mathcal{O}(\beta_{\rm QED}^{2}),$ (V.8)
with the leading angular-sector perturbation
$h_{\pm}(r)=4\,\bar{\Lambda}_{\pm}\,\frac{Q_{m}^{2}}{r^{4}},\qquad\text{where}\quad\Lambda_{\pm}=\beta_{\rm QED}\bar{\Lambda}_{\pm}\ \text{on the monopole background}.$ (V.9)
For $G=0$ and $L_{FG}=0$, Eq. (II.6) reduces to $\Lambda_{+}=L_{FF}$ and $\Lambda_{-}=\tfrac{1}{2}L_{GG}$, so that for EH $\bar{\Lambda}_{+}=8$ and $\bar{\Lambda}_{-}=7$ in the present conventions. (Any polarization-independent rescaling of $A_{\pm}$ and $B_{\pm}$ induced by $L_{F}$ is a common conformal factor and does not affect null geodesics; the birefringent input relevant to $\Delta\alpha$ is the mode dependence of $H_{\pm}$.)
@@ V.1.2 The finite-distance birefringence profile
To compute the differential bending angle $\Delta\alpha$, we utilize the decomposition into refractive curvature and path-induced contributions derived in Eq. (IV.8). The difference in the effective optical Gaussian curvature $\Delta K$ is the dominant source of birefringence. Given the $1/r^{4}$ scaling of the angular-sector perturbations $h_{\pm}$ in Eq. (V.9), the leading curvature difference scales as the second radial derivative of the corresponding metric deformation:
$\Delta K(r)\approx\frac{d^{2}}{dr^{2}}\!\left(\beta_{\rm QED}\,\Delta\bar{\Lambda}\,\frac{Q_{m}^{2}}{r^{4}}\right)\sim\beta_{\rm QED}\frac{Q_{m}^{2}}{r^{6}},$ (V.10)
where $\Delta\bar{\Lambda}\equiv\bar{\Lambda}_{+}-\bar{\Lambda}_{-}$. Inserting Eq. (V.10) into the refractive curvature integral of Eq. (IV.8), we integrate over the domain bounded by the photon path $r(\phi)\approx b/\sin\phi$. The area element contributes a factor of $\sqrt{\gamma}\sim r$, making the integrand scale as $\Delta(K\sqrt{\gamma})\sim r^{-5}$.
Performing the integration over the radial coordinate first, as prescribed in Eq. (IV.12), we find that the accumulated birefringence falls off rapidly with the impact parameter $b$. Specifically, the asymptotic birefringence scales as $\Delta\alpha_{\infty}\propto\beta_{\text{QED}}Q_{m}^{2}/b^{4}$. The steep fall-off ($n=4$) reflects the rapid decay of the monopole (spherically symmetric) vacuum-polarization contribution in the Euler–Heisenberg benchmark; by contrast, the dipole-like scaling discussed later corresponds to the steeper index $n=6$.
We can now explicitly write the finite-distance correction for QED vacuum birefringence. Substituting the scaling $n=4$ into the general series solution Eq. (IV.12), we obtain:
$\Delta\alpha_{\text{QED}}(r_{\text{R}},r_{\text{S}})=\Delta\alpha_{\infty}\left[1-c_{\text{R}}\left(\frac{b}{r_{\text{R}}}\right)^{4}-c_{\text{S}}\left(\frac{b}{r_{\text{S}}}\right)^{4}+\mathcal{O}\left(\frac{b^{6}}{r^{6}}\right)\right].$ (V.11)
@@ V.1.3 Physical implications
Equation (V.11) highlights a critical constraint for observational campaigns. The correction term scales as the fourth power of the ratio $b/r$. Unlike the gravitational deflection of light (which scales as $b/r$ in the finite-distance correction), the QED birefringence is highly localized near the source. Consequently, for an observer at infinity ($r_{\text{R}}\to\infty$), the finite-distance correction vanishes rapidly. However, if the source is located within the strong-field region (small $r_{\text{S}}$, e.g., emission from the surface of a magnetar), the truncation term $c_{\text{S}}(b/r_{\text{S}})^{4}$ becomes significant. The implication is that the standard asymptotic formulas over-estimate the expected polarization rotation for surface emission models, necessitating the use of Eq. (V.11) for precise polarimetric modeling of magnetars.
@@ V.2 Case Study 2: The Born-Infeld black hole
While Quantum Electrodynamics provides the standard birefringent vacuum response at low energies, Born-Infeld (BI) electrodynamics offers a classical non-linear extension of Maxwell’s theory, originally proposed to regularize the self-energy of point charges [49]. A key structural property of BI theory is that it is exceptional: the two polarization modes share the same characteristic cone (no vacuum birefringence) [13, 67]. We therefore use BI here as a consistency check of our finite-distance Gauss-Bonnet framework, emphasizing that our observable is the differential deflection $\Delta\alpha\equiv\alpha_{+}-\alpha_{-}$.
@@ V.2.1 Lagrangian and weak-field limit
The Born-Infeld Lagrangian density is given by
$L_{\rm BI}=\gamma^{2}\left[1-\sqrt{1+\frac{1}{2\gamma^{2}}F_{\mu\nu}F^{\mu\nu}-\frac{1}{16\gamma^{4}}\left(F_{\mu\nu}\tilde{F}^{\mu\nu}\right)^{2}}\right],$ (V.12)
@@
where $\gamma$ is the Born-Infeld coupling parameter (with dimensions of field strength). In the limit $\gamma\to\infty$ we recover Maxwell’s theory. To apply the weak-coupling expansion of Sec. IV we identify $\beta_{\rm BI}\propto\gamma^{-2}$.
@@
Expanding Eq. (V.12) for weak fields and retaining terms through ${\cal O}(\gamma^{-2})$, we obtain
$L_{\rm BI}\approx-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{1}{32\gamma^{2}}\Big[\big(F_{\mu\nu}F^{\mu\nu}\big)^{2}+\big(F_{\mu\nu}\tilde{F}^{\mu\nu}\big)^{2}\Big]+{\cal O}\!\left(\gamma^{-4}\right).$ (V.13)
Although $F_{\mu\nu}\tilde{F}^{\mu\nu}=0$ for purely electric or purely magnetic backgrounds, the full two-invariant structure of BI theory is essential for its characteristic (light-cone) geometry [13, 67].
V.2.2 Absence of birefringence
Born-Infeld theory is non-birefringent. That is, the Fresnel polynomial degenerates to a perfect square, so the two polarization modes propagate on the same effective null cone (up to conformal rescaling) [13, 67]. In our notation this implies
$g^{\mu\nu}_{{\rm eff},+}=g^{\mu\nu}_{{\rm eff},-},\qquad A_{+}=A_{-},\quad B_{+}=B_{-},\quad H_{+}=H_{-},$ (V.14)
and therefore the trajectory perturbations coincide, $\delta u_{+}=\delta u_{-}$. Substituting into the master formula Eq. (IV.8), both the refractive-curvature term and the path-induced term vanish, yielding
$\Delta\alpha_{\rm BI}(r_{R},r_{S})\equiv\alpha_{+}-\alpha_{-}=0.$ (V.15)
We stress that BI electrodynamics can still modify the common bending angle $\alpha_{+}=\alpha_{-}$ relative to Maxwell/GR, but this polarization-independent shift is not part of $\Delta\alpha$.
V.2.3 Multipole sensitivity of finite-distance suppression
The finite-distance correction series Eq. (IV.12) is sensitive to the multipole structure of the background field through the large-$r$ decay of $\Delta(K\sqrt{\gamma})$. For a dipole, the background field falls off as $\mathcal{B}\sim r^{-3}$, implying $F^{2}\sim r^{-6}$; correspondingly the leading birefringent curvature difference scales more steeply than in the monopole benchmark. Inserting this into Eq. (IV.8) yields the parametric scaling
$\delta\alpha_{\rm dipole}(r)\sim\int_{0}^{\arcsin(b/r)}(\sin\phi)^{6}\,d\phi\quad\Longrightarrow\quad n=6,$ (V.16)
demonstrating that the truncation terms $c_{S}(b/r_{S})^{n}$ can become even more important when higher multipoles contribute near the surface.